A sample of 20 pages was taken without replacement fromthe 1,591-page phone directory Ameritech Pages Plus Yellow pages
On each page, the mean area devoted to display ads was measured (a display as is a large block of multicolored illustrations, maps, and text). The data (in square millimeters) are shown below:
0; 260; 356; 403; 536; 0; 268; 369; 428; 536; 268;
396; 469; 536; 162; 338; 403; 536; 536; 130
a construct a 95% confidence interval for the true mean
b Why might normatlity be an issue here? c. What sample size would be needed to obtain an error of _+ 10 square millimeters with 99% confidence
d. If this is no a reasonable requirement, suggest one that is. (Data are from a project by MBA student Daniel R. Dalach)
I am new to this and have not been formally trained! Please help.

Answer by stanbon(75887)

A sample of 20 pages was taken without replacement fromthe 1,591-page phone directory Ameritech Pages Plus Yellow pages
On each page, the mean area devoted to display ads was measured (a display as is a large block of multicolored illustrations, maps, and text). The data (in square millimeters) are shown below:
0; 260; 356; 403; 536; 0; 268; 369; 428; 536; 268;
396; 469; 536; 162; 338; 403; 536; 536; 130
-----------------------------
a) construct a 95% confidence interval for the true mean
x-bar = 346.5
E = t*s/sqrt(n)
E = 2.093=79.738
95% CI: 346.5-79.738 < u < 346.5+79.738
95% CI: 266.762 < u < 426.24
----------------------------
b Why might normallity be an issue here?
All the pages were taken from the same phone book.
The CI is a statement about the whole population of phone books.
The sample might very well not be representative of the whole population.
------------------------------
c. What sample size would be needed to obtain an error of _+ 10 square millimeters with 99% confidence
Since E = t*s/(sqrt(n))
sqrt(n) = t*s/E
n = [t*s/E]^2 = [2.093*170.378/10]^2 = 1271.64
Rounded up you get n=1272
--------------------------
d. If this is no a reasonable requirement, suggest one that is.
You could lessen the sample size by decreasing the confidence
or increasing the error limit.
====================
Cheers,
Stan H.
----------------------------------